# Reducing GPU Memory Fragmentation via Spatio-Temporal Planning for Efficient Large-Scale Model Training

### v1

@@ 4. Allocation Profiler
Formally, we organize an allocation request and its associated free request into a memory request event $m$, which is defined as $m:=(s,t_{s},t_{e},p_{s},p_{e},d)$. Here, $s$ represents the request size; $t_{s}$ and $t_{e}$ are the allocation and free timestamps of the memory chunk, respectively; $p_{s}$ and $p_{e}$ identify the computation phases of allocation and free, respectively; and $d$ is a boolean flag indicating if the request originates from a dynamic layer (e.g., a MoE expert layer). For requests from dynamic layer (where $m.d=True$), two additional elements $l_{s}$ and $l_{e}$ are recorded, which are the originating module name for the allocation and free, respectively. This additional information allows us to group dynamic requests based on their temporal regularity, further details are provided in §5.2. Upon completion, the profiler outputs a list $\mathcal{M}$ of these characterized allocation requests, which is the primary input of the Plan Synthesizer.
@@ 5. Plan Synthesizer
The goal of the plan synthesizer is to produce a low fragmentation memory allocation plan that maximizes memory efficiency $E$ as defined in §2.2. Since allocated memory $M_{a}$ is fixed for a specific training configuration, the goal is then to minimize reserved memory $M_{r}$. To this end, the input of the synthesizer $\mathcal{M}$ is first partitioned into two subsets $\mathcal{M}_{s}:=\{m|m\in\mathcal{M},m.d=False\}$ and $\mathcal{M}_{d}:=\{m|m\in\mathcal{M},m.d=True\}$ according to their dynamicity. For $\mathcal{M}_{s}$ containing static request events, we perform static allocation planning to generate the Static Allocation Plan. Next, for dynamic request events $\mathcal{M}_{d}$, we find idle space in the plan (called Dynamic Reusable Space) that can be used to handle dynamic requests at runtime.
@@ 5.1. Static Allocation Planning
The Static Allocation Plan, denoted as $\mathcal{D}_{s}$, consists of a list of allocation decisions. Each allocation decision $d\in\mathcal{D}_{s}$ incorporates the six attributes of $m$ and is augmented with an additional attribute, $a$, which denotes the start address of the allocated memory chunk, i.e., $d:=m+(a)=(s,t_{s},t_{e},p_{s},p_{e},d,a)$. The allocation planning is then under the constraint that for any two allocation decisions $d_{i}$ and $d_{j}$, they cannot simultaneously have conflicting lifespans and conflicting address ranges. Otherwise, they will have intersecting memory and result in memory stomping.
@@
Since finding the optimal allocation plan is NP-hard and involves a large input scale as described in §1, brute-force methods or any pruning techniques that do not fundamentally reduce the complexity of the original search space (Maas et al., 2022; Moffitt, 2023) are infeasible. Inspired by the spatio-temporal regularity we uncover in §2.3, our idea is to decouple the searching in the space (i.e., memory address and size) and time (allocation time and free time) dimensions. At a high level, we first group memory allocation requests with the same temporal or spatial characteristics, devise a local plan for them to reduce the problem space, and then perform a global planning to reach the final plan. During the local and global planning, we greedily combine the optimal local memory layout for each request size to approximate the global optimal solution. We devise special abstractions for the temporal and spatial groups in the above algorithmic workflow, namely HomoPhase Group and HomoSize Group, respectively.
@@
HomoPhase Group Planning. A HomoPhase Group $\mathcal{M}_{g}$ contains allocation requests that start and end in the same computation phases, which is: $\mathcal{M}_{g}:=\{m\in\mathcal{M}_{s}\mid m.p_{s}=P_{s},\ m.p_{e}=P_{e}\}.$ Here, $P_{s}$ and $P_{e}$ denote a pair of computation phases (e.g., forward/backward passes), meaning all requests in $\mathcal{M}_{g}$ share similar lifespans.
@@
Since their lifespans overlap, packing them contiguously into a single memory block achieves local optimal. However, since their lifespans are only partially aligned, some memory may remain unused during parts of the timeline—these gaps are called spatio-temporal bubbles, causing memory fragmentation.
@@
To reduce such bubbles, we fuse adjacent groups when the end phase of one matches the start phase of another. The merged group can better reuse memory across phase boundaries. We evaluate memory efficiency using the time-memory product ($TMP$(Shore, 1975):
@@
As shown in Figure 6, we fuse two local plans $\mathcal{D}_{gi}$ and $\mathcal{D}_{gj}$ by inserting the smaller one into the larger. Assume $\mathcal{D}_{gi}.s>\mathcal{D}_{gj}.s$; we sort $\mathcal{D}_{gi}$ by end time in descending order and try placing each $d_{j}\in\mathcal{D}_{gj}$ at the lowest available address $addr$, starting from: $addr=\min_{d_{i}\in\mathcal{D}_{gi}}d_{i}.a.$ At each step:
1. Choose the earliest-starting $d_{j}$ that fits without conflict and place it at $addr$. Update $addr\leftarrow addr+d_{j}.s$.
2. If no fit is found, move $addr$ to the next $d_{i}.a$ in $\mathcal{D}_{gi}$:
@@
The fusion is accepted if the new $TMP$ improves over the weighted average of the originals, meaning fewer bubbles.
@@
HomoSize Group Planning. Allocation requests exhibit strong repetitiveness, with many requests having identical allocation sizes and differing only in their lifespan. This observation continues to hold true after HomoPhase Group planning, primarily because each microbatch exhibits identical behavior during training. Therefore, the HomoPhase Group formed through temporal grouping and fusion also possesses the characteristic that multiple such groups are identical in size, differing only in their lifespan.
Based on this observation, we propose a new abstraction termed HomoSize Group, which aggregates allocation requests of the same size property. For requests of a specific size $S$, there are only differences in their lifespan. Therefore, any subset of these requests with non-overlapping lifetimes can reuse the same space in GPU memory. In the time-space coordinate system, this shared space can be regarded as a layer within the memory space, referred to as a memory layer. To obtain a local optimal allocation plan for memory requests within a HomoSize Group, we need to minimize the number of memory layers required to allocate all requests.
Algorithm 1 describes the procedure of constructing memory layers for HomoSize Groups of a specific size $S$. To begin with, the allocation requests of size $S$ are included in a HomoSize Groups $\mathcal{M}_{s}$, and are sorted by their allocation time. Next, for each allocation request $m\in\mathcal{M}_{s}$, we try to find a memory layer whose last allocation request’s free time is closest to but smaller than $m$’s allocation time, so as to minimize the idle time of the layer while avoiding conflicting lifespans. If we can find such a layer, $m$ is appended to the layer’s tail. Otherwise, a new layer is constructed and is populated by $m$. In this way, we minimize both the intra-layer gaps and the total number of layers.
Global Planning. In the local planning, memory requests are grouped based on their temporal and spatial regularities into HomoPhase Groups and HomoSize Groups, and generate a local memory allocation plan for each group. These local plans can then be regarded as building blocks of the global memory allocation plan, as shown in Figure 6.
Specifically, the first step in generating the global allocation plan is to partition all memory requests within one training iteration into different HomoPhase Groups based on their temporal characteristics. Adjacent HomoPhase Groups are then merged to produce local plans, denoted as $\mathcal{D}_{g}$. These local plans are subsequently treated as unified memory requests in the next stage of spatial grouping, where they are classified into different HomoSize Groups according to their allocation sizes. Each HomoSize Group then constructs memory layers within the group. We execute the construction in descending order of the request size within the HomoSize Group, since smaller memory requests may fit into the unused intervals of memory layers allocated for larger requests, thus improving overall memory efficiency. Therefore, before constructing memory layers for size $S_{i}$ (at which point all larger groups have already been processed), we first attempt to insert the requests of size $S_{i}$ into the free intervals of existing larger memory layers. Any remaining requests of size $S_{i}$ that cannot be placed in the larger layers are then assigned new memory layers by executing the Algorithm 1. Finally, after all layers have been constructed, each memory request is assigned a specific address within the global allocation plan.
@@
Output : MemoryLayer List $\mathcal{L}=(\mathcal{M}_{l1},\mathcal{M}_{l2},...)$
1 $\mathcal{L}\leftarrow[\>]$;
2 $\mathcal{M}_{s}$.sort(key=$m.t_{s}$);
@@ 1 $\mathcal{L}\leftarrow[\>]$;
6             $\mathcal{M}_{l}\leftarrow$ new MemoryLayer(size=$m.s$);
@@
Algorithm 1 Memory Layer Construction for HomoSize Group.
@@ 5.2. Locating Dynamic Reusable Space
Dynamic allocation requests ($\mathcal{M}_{d}$) are characterized by sizes determined only at runtime, necessitating online allocation. A key insight from our profiler analysis is that the peak memory usages of static and dynamic allocations typically do not occur simultaneously. Therefore, managing static and dynamic allocations in separate regions—where each region is provisioned for its respective peak—results in memory fragmentation. For instance, when dynamic memory usage peaks, the dedicated static memory region often contains substantial idle capacity. This observation of underutilized memory, a direct consequence of misaligned peak demands under segregated management, forms the primary motivation for our proposed dynamic planning strategy.
@@
Leveraging the predictable lifetimes of dynamic memory allocations, our approach proactively identifies reusable memory regions within the Static Allocation Plan before runtime. To achieve fine-grained temporal precision for these dynamic requests—in contrast to the computation-phase granularity used for static allocations—we operate at the model layer level. This refined granularity enables a more precise interrogation of Dynamic Reusable Space. Such accurately identified regions within the typically shorter layer-level intervals tend to be more effectively utilized, maximizing opportunities for dynamic memory reuse and thereby reducing the peak GPU memory footprint. We characterize each dynamic request by its malloc layer, $l_{s}$, and its free layer, $l_{e}$ (profiling methodology in §4). This ($l_{s}$, $l_{e}$) pair establishes a bounding temporal interval, from $l_{s}$’s start to $l_{e}$’s end, which contains the lifespan of the dynamic allocation. To systematically manage these lifetimes, we classify all dynamic allocation requests into distinct groups, called HomoLayer Group, where each group $\mathcal{G}$ comprises requests sharing identical ($l_{s}$, $l_{e}$) pairs:
@@
The identified Dynamic Reusable Space $\mathcal{A}_{i}$ are subsequently designated as candidate reusable regions. At runtime, when a dynamic allocation request belonging to a particular ($l_{s}$,$l_{e}$) group arises, the allocator can preferentially utilize these pre-vetted regions, thereby ensuring that dynamic allocations are placed in memory spaces that will not conflict with future, planned static allocations.
@@ 6. Runtime Allocation
The runtime allocator manages the GPU memory and serves allocation requests based on the allocation plan generated by the plan synthesizer. It consists of two main components, a static allocator that handles allocation requests without runtime dynamics (i.e., $m.d==False$), and a dynamic allocator that handles allocations with runtime dynamics (i.e., $m.d==True$). During runtime, when an allocation request is received by the Runtime Allocator, the Request Matcher routes the request to an appropriate allocator based on the dynamic characteristics of the current model layer (detail shows in §7). Furthermore, to address scenarios such as potential mismatch between actual runtime allocation requests and the Static Allocation Plan, or instances of inadequate Dynamic Reusable Space for dynamic requests, STWeaver’s runtime allocation further incorporates a caching allocator. This component is designed to manage these exceptional cases, thereby guaranteeing the overall robustness of the system.
@@ 6.2. Dynamic Allocator
The primary strategy of the dynamic allocator is to prioritize reusing memory from the static memory pool, which was pre-allocated for predictable requests. To prevent conflicts, STWeaver meticulously tracks all currently available address intervals ($\mathcal{A}_{a}$) in this pool. When any memory block is allocated or freed, $\mathcal{A}_{a}$ is updated accordingly.
When a dynamic request $m$ arrives, the allocation process begins by first identifying the Dynamic Reusable Space $\mathcal{A}_{i}$, which is the available space in static memory pool for the HomoLayer Group contains $m$. Since prior allocations may have already occupied parts of this space, we must identify the portions that are still free. To find the actual memory available for allocation, STWeaver computes the intersection of this potential space $\mathcal{A}_{i}$ with the currently free intervals $\mathcal{A}_{a}$. This calculation yields a set of candidate intervals, $\mathcal{A}_{c}(m)$:
@@
If no candidate interval in the static pool can satisfy the request, the system falls back to the caching allocator as a secondary option.

### v2

@@ 4. Allocation Profiler
Formally, we organize an allocation request and its associated free request into a memory request event $m$, which is defined as $m:=(s,t_{s},t_{e},p_{s},p_{e},dyn)$. Here, $s$ represents the request size; $t_{s}$ and $t_{e}$ are the allocation and free timestamps of the memory chunk, respectively; $p_{s}$ and $p_{e}$ identify the computation phases of allocation and free, respectively; and dyndyn is a boolean flag indicating if the request originates from a dynamic layer (e.g., a MoE expert layer). For requests from dynamic layers (where $m.dyn=True$), two additional elements $l_{s}$ and $l_{e}$ are recorded, which are the originating module name for the allocation and free, respectively. This additional information allows us to group dynamic requests based on their temporal regularity, further details are provided in §5.2. Upon completion, the profiler outputs a list $\mathcal{M}$ of these characterized allocation requests, which is the primary input of the Plan Synthesizer.
@@ 5. Plan Synthesizer
The goal of the plan synthesizer is to produce a low fragmentation memory allocation plan that maximizes memory efficiency $E$ as defined in §2.2. Since allocated memory $M_{a}$ is fixed for a specific training configuration, the goal is then to minimize reserved memory $M_{r}$. To this end, the input of the synthesizer $\mathcal{M}$ is first partitioned into two subsets $\mathcal{M}_{s}:=\{m|m\in\mathcal{M},m.dyn=False\}$ and $\mathcal{M}_{d}:=\{m|m\in\mathcal{M},m.dyn=True\}$ according to their dynamicity. For $\mathcal{M}_{s}$ containing static request events, we perform static allocation planning to generate the Static Allocation Plan. Next, for dynamic request events $\mathcal{M}_{d}$, we find idle space in the plan (called Dynamic Reusable Space) that can be used to handle dynamic requests at runtime.
@@ 5.1. Static Allocation Planning
The Static Allocation Plan, denoted as $\mathcal{D}_{s}$, consists of a list of allocation decisions. Each allocation decision $d\in\mathcal{D}_{s}$ incorporates the six attributes of $m$ and is augmented with an additional attribute, $a$, which denotes the start address of the allocated memory chunk, i.e., $d:=m+(a)=(s,t_{s},t_{e},p_{s},p_{e},dyn,a)$. The allocation planning is then under the constraint that for any two allocation decisions $d_{i}$ and $d_{j}$, they cannot simultaneously have conflicting lifespans and conflicting address ranges. Otherwise, they will have intersecting memory and result in memory stomping.
@@
Since finding the optimal allocation plan is NP-hard and involves a large input scale as described in §1, brute-force methods or any pruning techniques that do not fundamentally reduce the complexity of the original search space (telamalloc; minimalloc) are infeasible. Inspired by the spatio-temporal regularity we uncover in §2.3, our idea is to decouple the searching in the space (i.e., memory address and size) and time (allocation time and free time) dimensions. In the global planning process, requests that exhibit temporal or spatial regularities are handled through local planning, where efficient layouts are derived by exploiting these regularities. The resulting local plans then become integral components of the global allocation plan. To support this workflow, we introduce two abstractions: HomoPhase Group and HomoSize Group. The HomoPhase Groups gather allocation requests with similar lifespan, and the HomoSize Group gather requests with same size. We will further introduce them in the following.
Global Planning.   At a high level, the components of the global planning are derived through local plans. Each local plan groups allocation requests that exhibit regularity along one dimension (spatial or temporal), allowing us to exploit such regularities to design allocation algorithms that improve memory efficiency.
Specifically, the first step in generating the global allocation plan is to partition all memory requests within one training iteration into different HomoPhase Groups based on their temporal characteristics. Adjacent HomoPhase Groups are then merged to produce local plans, denoted as $\mathcal{D}_{g}$ (see in HomoPhase Group Planning, Figure 6 upper left). These local plans are subsequently treated as unified memory requests in the next stage of spatial grouping, where they are classified into different HomoSize Groups according to their allocation sizes. Each HomoSize Group will construct its own local plan. We execute the construction in descending order of the request size for the HomoSize Group, since smaller memory requests may fit into the unused intervals of larger requests, thus improving overall memory efficiency. Before planning for a HomoSize Group of size $S_{i}$, all larger groups have already been processed. At this stage, we first try to place the requests of size $S_{i}$ into the free intervals of the existing local plans of larger groups. The remaining requests of size $S_{i}$ that cannot be placed in the larger group plans will then construct a new local plan (see in HomoSize Group Planning, Figure 6 bottom left). Finally, after all local plans for HomoSize Group have been constructed, each memory request is assigned a specific address within the global allocation plan (Figure 6 right).
Next, we introduce how local plans for HomoPhase Group and HomoSize Group are generated.
Figure 7. Examples of fusion between different HomoPhase Groups. In the left example, the fusion leads to an increase in TMPTMP (‘WA’ means weighted average), so the fusion scheme is adopted; in the right example, the fusion results in a decrease in TMPTMP, so it is rejected.
@@
HomoPhase Group Planning.   A HomoPhase Group $\mathcal{M}_{g}$ contains allocation requests that start and end in the same computation phases, which is: $\mathcal{M}_{g}:=\{m\in\mathcal{M}_{s}\mid m.p_{s}=P_{s},\ m.p_{e}=P_{e}\}.$ Here, $P_{s}$ and $P_{e}$ denote a pair of computation phases (e.g., forward/backward passes), meaning all requests in $\mathcal{M}_{g}$ share similar lifespans. $\mathcal{D}_{g}$ is the allocation plan for HomoPhase Group $\mathcal{M}_{g}$, where each allocation request are placed with a relative address.
@@
Since their lifespans overlap, packing them contiguously into a single memory block achieves local optimal. However, their lifespans are only partially aligned, some memory may remain unused during parts of the timeline. These gaps are called spatio-temporal bubbles, causing memory fragmentation.
@@
To reduce such bubbles, we fuse adjacent groups when the end phase of one matches the start phase of another. The merged group can better reuse memory across phase boundaries. We evaluate memory efficiency using the time-memory product (TMPTMP(best-fit-allocation):
@@
As shown in Figure 6 upper left, we fuse two local plans $\mathcal{D}_{gi}$ and $\mathcal{D}_{gj}$ by inserting the smaller one into the larger. Assume $\mathcal{D}_{gi}.s>\mathcal{D}_{gj}.s$; we sort $\mathcal{D}_{gi}$ by end time in descending order and try placing each $d_{j}\in\mathcal{D}_{gj}$ at the lowest available address addraddr, starting from: $addr=\min_{d_{i}\in\mathcal{D}_{gi}}d_{i}.a.$ At each step:
1. Choose the earliest-starting $d_{j}$ that fits without conflict and place it at addraddr. Update $addr\leftarrow addr+d_{j}.s$.
2. If no fit is found, move addraddr to the next $d_{i}.a$ in $\mathcal{D}_{gi}$:
@@
The fusion is accepted only if the new $TMP$ increases over the weighted average of the originals, meaning fewer bubbles, as shown in Figure 7.
@@
HomoSize Group Planning.
@@
[…]
1
21ex[$\mathcal{L}\leftarrow$[\>];
[…]
@@
7     $\mathcal{M}_{l}\leftarrow$ new Memory-Layer(size=$m.s$);
@@
Algorithm 1 Memory-Layer Construction for HomoSize Group.
Allocation requests exhibit strong repetitiveness, with many requests having identical allocation sizes and differing only in their lifespan. This observation continues to hold true after HomoPhase Group planning, primarily because each microbatch exhibits identical behavior during training. Therefore, the HomoPhase Group formed through temporal grouping and fusion also possesses the characteristic that multiple such groups are identical in size, differing only in their lifespan.
Based on this observation, we propose the abstraction termed HomoSize Group, which aggregates allocation requests of the same size property. For requests of a specific size $S$, there are only differences in their lifespan. Therefore, any subset of these requests with non-overlapping lifetimes can reuse the same space in GPU memory. In the time-space coordinate system, this shared space can be regarded as a layer within the memory space, referred to as a memory-layer. To obtain a local optimal allocation plan for memory requests within a HomoSize Group, we need to minimize the number of memory-layers required to allocate all requests.
Algorithm 1 describes the procedure of constructing memory-layers for HomoSize Groups of a specific size $S$. To begin with, the allocation requests of size $S$ are included in a HomoSize Groups $\mathcal{M}_{s}$, and are sorted by their allocation time (Line 1). Next, for each allocation request $m\in\mathcal{M}_{s}$, we try to find a memory-layer whose last allocation request’s free time is closest to but smaller than $m$’s allocation time, so as to minimize the idle time of the memory-layer while avoiding conflicting lifespans (Line 1). If we can find such a memory-layer, $m$ is appended to the layer’s tail. Otherwise, a new memory-layer is constructed and is populated by $m$ (Line 1 - 1). In this way, we minimize both the intra-layer gaps and the total number of memory-layers.
@@ 5.2. Locating Dynamic Reusable Space
Dynamic allocation requests ($\mathcal{M}_{d}$) are characterized by sizes determined only at runtime, necessitating online allocation. A key insight from our profiler analysis is that the peak memory usages of static and dynamic allocations typically do not occur simultaneously. When static and dynamic allocations are managed in separate memory regions, each region must be provisioned for its individual peak usage. However, since the peaks of static and dynamic demands occur at different times, the reserved capacity of one region (e.g., static) often remains idle while the other (e.g., dynamic) reaches its peak. This temporal mismatch results in significant underutilization of memory resources and lowers overall efficiency. To overcome this limitation, our dynamic planning strategy reuses the idle space of static allocations to accommodate dynamic requests.
@@
Leveraging the predictable lifetimes of dynamic memory allocations, our approach proactively identifies reusable memory regions within the Static Allocation Plan before runtime. In contrast to the computation-phase granularity used for static allocations, here we operate at the model layer level to achieve fine-grained temporal precision for these dynamic requests. This refined granularity enables a more precise interrogation of Dynamic Reusable Space. Since these regions lie within shorter layer-level intervals, they can be utilized more effectively, creating more opportunities for dynamic memory reuse and thereby lowering the peak GPU memory footprint. We characterize each dynamic request by its malloc model layer, $l_{s}$, and its free model layer, $l_{e}$ (profiling methodology in §4). This ($l_{s}$, $l_{e}$) pair establishes a bounding temporal interval, from $l_{s}$’s start to $l_{e}$’s end, which contains the lifespan of the dynamic allocation. To systematically manage these lifetimes, we classify all dynamic allocation requests into distinct groups, called HomoLayer Group, where each group $\mathcal{G}$ comprises requests sharing identical ($l_{s}$, $l_{e}$) pairs:
@@
The identified Dynamic Reusable Space $\mathcal{A}_{i}$ are subsequently designated as candidate reusable regions. At runtime, when a dynamic allocation request belonging to a particular ($l_{s}$$l_{e}$) group arises, the allocator can preferentially utilize these pre-vetted regions, thereby ensuring that dynamic allocations are placed in memory spaces that will not conflict with future, planned static allocations.
@@ 6. Runtime Allocation
The runtime allocator manages the GPU memory and serves allocation requests based on the allocation plan generated by the plan synthesizer. It consists of two main components, a static allocator that handles allocation requests without runtime dynamics (i.e., $m.dyn==False$), and a dynamic allocator that handles allocations with runtime dynamics (i.e., $m.dyn==True$). During runtime, when an allocation request is received by the Runtime Allocator, the Request Matcher routes the request to an appropriate allocator based on the dynamic characteristics of the current model layer (detail shows in §8). Furthermore, to address scenarios such as potential mismatch between actual runtime allocation requests and the Static Allocation Plan, or instances of inadequate Dynamic Reusable Space for dynamic requests, STAlloc’s runtime allocation further incorporates a caching allocator. This component is designed to manage these exceptional cases, thereby guaranteeing the overall robustness of the system.
@@ 6.2. Dynamic Allocator
The primary strategy of the dynamic allocator is to prioritize reusing memory from the static memory pool, which was pre-allocated for predictable requests. To prevent conflicts, STAlloc meticulously tracks all currently available address intervals ($\mathcal{A}_{a}$) in this pool. When any memory block is allocated or freed, $\mathcal{A}_{a}$ is updated accordingly.
When a dynamic request $m$ arrives, the allocation process begins by first identifying the Dynamic Reusable Space $\mathcal{A}_{i}$, which is the available space in static memory pool for the HomoLayer Group contains $m$. Since prior allocations may have already occupied parts of this space, we must identify the portions that are still free. To find the actual memory available for allocation, STAlloc computes the intersection of this potential space $\mathcal{A}_{i}$ with the currently free intervals $\mathcal{A}_{a}$. This calculation yields a set of candidate intervals, $\mathcal{A}_{c}(m)$:
@@
If no candidate interval in the static pool can satisfy the request, the system falls back to the caching allocator as a secondary option. This caching allocator follows PyTorch’s CUDA Caching Allocator, reusing previously allocated blocks through a block management mechanism.
7. Complexity Analysis
We analyze the computational complexity of our method. The discussion is organized into two main parts: (i) plan synthesis, which is performed once before execution, and (ii) runtime allocation, which is invoked during model training or inference.
7.1. Plan Synthesis
Static Requests. Plan synthesis for static requests consists of two stages. First, requests are grouped into HomoPhase Groups and fused when possible, which requires sorting by request endpoints and costs $O(N\log m)$, where $N$ is the total number of requests and $m$ is the maximum group size. Second, the resulting groups are ordered and inserted by size. Sorting by size requires $O(N\log N)$, while the layer construction within each group (Algorithm 1) is linear in group size. Thus, static plan synthesis has overall complexity $O(N\log N)$.
Dynamic Requests. For dynamic requests, all $k$ HomoLayer Groups’ temporal intervals are known in advance. By sorting time intervals and performing a batched sweep, the complexity is:
$O\big(N\log N+k\log(N+k)+\sum_{i=1}^{k}r_{i}\big),$
where $r_{i}$ is the number of static requests overlapping the $i$-th query. Since typically $k\ll N$ and $\sum_{i}r_{i}\ll kN$, the dynamic part is asymptotically bounded by $O(N\log N)$.
7.2. Runtime Allocation
Static Requests. At runtime, static requests incur $O(1)$ lookup cost because their addresses are pre-computed in the plan.
Dynamic Requests. Dynamic requests are allocated by intersecting pre-computed reusable spaces with the currently active blocks. If $n$ is the number of active blocks at that moment, this step costs $O(n)$. In practice $n$ is orders of magnitude smaller than $N$, ensuring that runtime allocation remains efficient.
Overall, plan synthesis is dominated by $O(N\log N)$ preprocessing cost, while runtime allocation requires only $O(1)$ per static request and $O(n)$ per dynamic request. This guarantees that both initialization and runtime operations are efficient in large-scale training.